OpenStudy (anonymous):

If $500 is invested at 8.5% simple interest for 6 years, what will be the total amount of money in the investment account at the end of the 6 years?

OpenStudy (anonymous):

I=Prt = 500 *0.085*6=255 ===> total amount = 500 +255=$755